What are three quadratic equations with a,b and c (coefficients of x^2,x and the constant.
using
1. Integers
2. Rational numbers
3. Irrational numbers

Answer by stanbon(75887)

What are three quadratic equations with a,b and c (coefficients of x^2,x and the constant.
using
1. Integers:
(x-2)(x+5) = x^2 +3x -10 has integral solutions.
-----------------------------
2. Rational numbers
(2x-1)(3x-2) = 6x^2 - 7x +6 has rational solutions
-----------------------------
3. Irrational numbers
(x-sqrt(2))(x+sqrt(3)) = x^2 + [sqrt(3)-sqrt(2)]x - sqrt(6) has
irrational solutions.
==========================
Cheers,
Stan H.